OpenStudy (jackfrost33):

HELP!!! PLEASE how would you solve this with out getting a decimal i have tried but no answer i got was right, 5u+2v=18 and 6u+v=22 .

zepdrix (zepdrix):

What do you mean "without getting a decimal"? If a decimal value is the answer, then that's the answer... Would you rather write it as a fraction or something?

OpenStudy (jackfrost33):

my teacher told us its not going to be a decimal answer

zepdrix (zepdrix):

Oh I see :)

zepdrix (zepdrix):

\[\large\rm 5u+2v=18\]\[\large\rm 6u+~~v=22\]Personally I really like the elimination method: We try to match up one of the variables in each equation, and then subtract.

OpenStudy (jackfrost33):

yeah... its to be done using substuiton method but i dont really get that either

zepdrix (zepdrix):

Substitution? Fine fine fine :p

zepdrix (zepdrix):

Notice the v in the second equation has no fancy coefficient. So it will be easy to write it in terms of v=(stuff)

zepdrix (zepdrix):

\[\large\rm 6u+v=22\]How bout we subtract 6u from each side, ya?

OpenStudy (jackfrost33):

yes i am

OpenStudy (jackfrost33):

but the problem it 22 not 23

zepdrix (zepdrix):

This system, the way you wrote it, has ugly decimal values for a solution.

OpenStudy (sooobored):

23 would make it a nice problem

OpenStudy (jackfrost33):

ohhhkay ill tell my teacher in the morning about thais maybe i wont have to do this

OpenStudy (jackfrost33):

ohhhkay ill tell my teacher in the morning about thais maybe i wont have to do this

zepdrix (zepdrix):

XD

OpenStudy (sooobored):

well, i still suggest doing it if it were 23

OpenStudy (sooobored):

good practice

OpenStudy (jackfrost33):

okay ill try it with 23

zepdrix (zepdrix):

\[\large\rm 6u+v=23\]In the second equation, subtracting 6u from each side gives us,\[\large\rm v=23-6u\]understand that step?

OpenStudy (jackfrost33):

yeah

zepdrix (zepdrix):

This equality sign is telling us that \(\large\rm \color{orangered}{v}\) is the same as \(\large\rm \color{orangered}{(23-6u)}\). \[\large\rm 5u+2\color{orangered}{v}=18\] So do you understand how we can use this with our other equation?

OpenStudy (jackfrost33):

not really but i think i can do that

zepdrix (zepdrix):

We can replace the v in our first equation with 23-6u because those things are equivalent,\[\large\rm 5u+2\color{orangered}{(23-6u)}=18\]

zepdrix (zepdrix):

From that point, you have an equation involving only u, no more v, so you can do some quick steps and solve for u.

OpenStudy (jackfrost33):

okay

OpenStudy (seratul):

Here is a simpler way In the beginning you said that we can rewrite the equation to (v=23-6u) We can also rewrite 5u+2v=18 to 2v=18-5u So we should now have the two equation of v=23-6u and 2v-18-5u This means that 2(23-6u)=18-5u It can simplify to 46-12u=18-5u And then you just solve. 46=18+7u 28=7u u=4 Then just plug it into the equation to find v. You don't have to do it this way, but I just really wanted to say it :D It's good review for SAT.

OpenStudy (jackfrost33):

thanksguys !!!!:)